Two bikers started traveling at the same corner, one going south, the other going west. One biker is traveling at 3 km/h faster than the other, After one hour, the two bikers are 15 kilometers apart. Find the rate of each.
Thanks!

Answer by ankor@dixie-net.com(22740)

Two bikers started traveling at the same corner, one going south, the other going west.
One biker is traveling at 3 km/h faster than the other,
After one hour, the two bikers are 15 kilometers apart.
Find the rate of each.
:
Let s = the speed of the slower biker
then
(s+3) = speed to the faster
dist = speed * time, therefore:
1s = distance traveled by the slower in 1 hr
1(s+3) = distance traveled by the faster
:
this is a pythag problem a^2 + b^2 = c^2, in this problem
a = 1s
b = 1(s+3)
c = 15
:
we don't need the one, we have
s^2 + (s+3)^2 = 15^2
:
FOIL(s+3)(s+3)
s^2 + s^2 + 6s+ 9 = 225
:
combine on the left as a quadratic equation
2s^2 + 6s + 9 - 225 = 0
2s^2 + 6s - 216 = 0
:
simplify, divide by 2
s^2 + 3s - 108 = 0
:
Factors to
(s+12)(s-9) = 0
:
the positive solution is all we want here
s = 9 mph is the speed of the slower bike
:
You can find the rate of the faster one, check it in the pythag equation